# Motion simulation of radio-labeled cells in whole-body positron emission tomography

### v2

@@ 2.4 PET simulation
Raw PET data is generated using the Geant4 application for tomographic emission GATE (version 9.0) [39, 18]. The cell paths outputs from the CeFloPS simulations have the format of GATE placement files and thus are used as generic moving sources in the simulations. The sources, i.e. the cells, are simulated as spheres whose diameters and activities can be specified by the user. For the example shown in this work, a diameter of 10 µm and an activity of 1 kBq per cell were simulated for 100 cells. This activity per cell and number of cells does not correspond to a realistic scenario and was only chosen to serve as a proof-of-principle. The PET scanner used in the simulation is based on the Siemens Biograph Vision Quadra (Siemens Healthineers, Erlangen, Germany), a large-body PET system using approximately 250000 single lutetium oxyorthosilicate (LSO) detector elements covering an axial field of view (FOV) of 106 cm [40]. The background radiation in the scanner generated by the presence of 176Lu in the LSO crystals can be simulated if desired but was neglected in the current simulation. The attenuation map for gamma energies of 511 keV generated from the XCAT phantom was used as an attenuation phantom. The attenuation phantom was placed in the PET scanner so that the upper body including the head were in the FOV of the scanner. Since the simulated cell paths are based on the blood vessel and organ structure of the XCAT phantom, their positions correspond exactly to the anatomical structures in the attenuation phantom, making the simulation also realistic in view of the proportion of attenuated and scattered events. Although the lower body exceeds the FOV of the scanner, it is still simulated in case a cell is located there.
@@
Starting at the aorta with a mean blood flow of 100100100~{}100ml/s, the blood distributes into all vessels and tissues depending on the resistances of the vessel system and the blood pressure differences (as listed exemplarily in Table 2). It is evident that the lungs receive all blood through the contraction of the right myocardium, which accounts for 100 % of the total blood flow. Notably, the kidneys, gastrointestinal tract, and liver show strong alignment between the simulated flow rate and physiological flow distributions from literature [41], highlighting the accuracy of CeFLoPS in capturing predominant systemic blood flow patterns. However, considerable discrepancies are observed for the brain, spleen, pancreas and thyroid. This behavior can be attributed to the insufficient definition of the vascular structure of these organs or a high vessel resistance due to a large number of small vessels in the case of the brain.
Table 2: Blood flow distribution calculated by CeFloPS
@@
(of total blood flowa)
Typical blood flow rate
in humanb
(of total blood flow)
@@
Fig. 6 depicts a maximum intensity projection (MIP) of the distribution of 100,000 injected cells over the first 100 seconds following injection into the left arm vein. As would be anticipated, the route from the injection site to the heart, the heart itself, the pulmonary arteries and veins, the aorta and arteries close to the aorta demonstrate a high degree of frequency. In regions such as the legs and arms, it is evident that certain arteries and veins have yet to be reached by a significant number of cells. Additionally, there are observable transitions of cells into the organs and reflux into the veins, particularly evident in the kidneys, liver, and vena cava.
Figure 6: Maximum intensity projection of the the vessels and tissues traversed by 100,000 cells injected into the right arm vein during the first 100 seconds of a simulation. Darker vessels/tissues are traversed by more cells.
The biased 3D random walk that cells perform when they migrate from the blood into a tissue is illustrated in Fig. 7a using the example of the lung for a duration of 2000 s. Simulation outcome is visualized through MIPs in three orthogonal planes to illustrate spatial traversal patterns clearly. The distribution of cells entering the random walk through the arterial sources is weighted according to simulated arterial blood flow fractions of the vessels in the lung. It prominently demonstrates regions of enhanced cellular traversal intensity, indicative of high-flow vascular pathways, and shows that the biased random walk ensures that the entire organ tissue is potentially accessible by incoming cells. Furthermore, the temporal dynamics of cell migration are elucidated in Figure Fig. 7b, where the histogram of random walk durations reveals a rapid traversal of the lung vasculature for the majority of cells. The mean duration is 3.77 s, which is consistent with the theoretical specification outlined in [30].
@@
Fig. 8 illustrates the locations of 100 simulated cells within a node system at six time points, compared to MIPs reconstructed from PET data using an expectation maximization (EM) algorithm without attenuation correction. The reconstructed positions and cell counts align well with ground-truth positions, despite minor reconstruction artifacts in the MIPs causing low-intensity signals outside the body due to missing attenuation and scatter correction. Initially, cells (red spheres) are injected into the left arm vein (median cubital) over three seconds. After six seconds, cells migrate towards the left clavicle, appearing as a single dark spot in the MIP, making individual cells indistinguishable. By 18 seconds, most cells distribute within the lung, with some transitioning to arterial blood, allowing individual cells and clusters to be distinguishable in the MIP. After 30 seconds, cells predominantly remain in the lung; some transition to compartments $C_{1}$ or $C_{2}$, changing color from red to green, while others disperse through the abdominal cavity and neck, identifiable in the MIP. At one minute, cells migrate further into the torso, with some returning to venous blood and lungs, and individual cells reach the head, right arm, and left leg (outside the scanner’s FOV). After five minutes, cells disperse body-wide, with some transitioning to compartment $C_{1}$ and $C_{2}$, as seen in cells in the right arm; one cell remains, while the other leaves the arm. This dispersion continues over ten minutes, increasing the proportion of stationary cells in compartment $C_{1}$ and $C_{2}$, thus facilitating the identification of individual cells in MIPs over time due to wider distribution and fewer clusters.
@@
Figure 8: Comparison of simulated ground-truth cell positions and reconstructed PET data for 100 cells at six different time points showing the positions of both moving cells (red, left) in the node system as well as stationary cells that have entered $C_{1}$ or $C_{2}$ (green, left). The reconstructed PET maximum intensity projections (right) show the cell distribution (black) superimposed to the organ and blood vessel structure. The red dotted line marks the FOV of the simulated PET scanner.
@@
Figs. 9 and 10 illustrate the temporal distribution of 100,000 cells in arterial blood, lung, kidneys, liver, spleen, and pancreas as absolute cell numbers (sum of cells in $C_{1}$ and $C_{2}$) and as cell concentrations per milliliter tissue, respectively. Initially, the cell count in the arterial blood rapidly drops and a pronounced peak of cell presence is observed within lung, liver, and kidneys, indicative of early vascular dissemination and transitions from $C_{A}$ to $C_{1}$. Subsequently, cell counts decline within these tissues, aligning with physiological expectations of cell transitioning back from $C_{1}$ to $C_{A}$ and leaving the tissue into the venous blood. Notably, over extended simulation time, the cell count in the kidneys rises again and the kidneys and the pancreas exhibit a clear progressive increase in cell accumulation, indicating trapping of cells in $C_{2}$. Meanwhile, the spleen demonstrates a relatively steady cell population, reflecting a possible equilibrium between cell influx and clearance. The concentration plots (Fig. 10) underscore these observations by correcting for organ volume, thus offering additional insight into organ-specific cell affinity beyond raw cell counts. In terms of raw cell count, spleen and pancreas reach significantly lower numbers than the other tissues, which is not noticeable in the concentration plot. Notably, the kidneys exhibit considerably higher values than other tissues with regard to both the absolute number of cells and the cell concentration.
@@
Fig. 11 illustrates the temporal evolution of the 100,000 simulated cells across the virtual, blood, and tissue compartments depicted in Fig. 4 and described by (11). After the 30 second long injection process, all cells have transitioned from $C_{\mathrm{inj}}$ to $C_{A}$. This is followed by a swift redistribution into tissues or removal from blood and transition into $C_{\mathrm{rem}}$. After a brief transient phase, a stable equilibrium emerges, characterized predominantly by cells in the removed state with over 80%. The dynamics clearly show that only a small fraction of injected cells persist within blood or the defined tissue compartments at longer time points. The majority of cells is located in tissues without specified dynamics, like skin, muscle or adipose.
@@
Figure 11: Temporal distribution of 100,000 simulated cells across compartments shown in Fig. 4: injected, arterial blood, tissue, and removed).
@@
Fig. 12 compares the simulated cell concentrations in the tissues to the the distributions one would expect from (10). To calculate the theoretically expected distributions for each tissue, the number of cells perfusing from the arterial blood into tissue are mapped. The arterial input to each tissue is determined by the unique and locally dependent distribution of blood flow and the subsequent delay and dispersion effects. Together with the transition rates (Table LABEL:tab:k_values) of each tissue, the arterial input allows to derive the expected cell distributions from the differential equations. The reference volume to which the cells of the arterial input function are normalized in order to obtain a concentration is assumed to be a tissue-specific variable. This variable is dependent on the volume of the small arteries, arterioles and capillaries leading to and perfusing the tissue. This factor is determined by fitting the two distributions and multiplying the result with the arterial input function. This yields the concentration of cells in the arterial blood and the expected theoretical curve. As seen in fig. 12 the simulated distributions describe the theoretically expected ones well, with small over- or undershoots of the initial peaks. In the case of the spleen and pancreas, the theoretical expected cell concentration exhibits various fluctuations, due to the small number of cells in these tissues.
@@
Figure 12: Simulated concentration distributions profiles of cells in tissues and tissue specific arterial blood input. Solid colored lines represent the simulated cell concentrations per milliliter in the respective tissues, while the red lines show the cell concentration in the arterial blood perfusing each tissue. The plot in the upper left displays the total arterial blood concentration in the whole body. The dashed black lines depict the theoretically expected concentration distributions derived from (10), based on the tissue specific arterial input functions and transition rates (Table LABEL:tab:k_values).
The performance of the simulation is linked to the parameters chosen to prepare the data as well as parameters regarding a desired output. We make use of multiprocessing with python and launch multiple instances that each simulate a fraction of all cells. By default, no synchronization between those processes is necessary, as only positional data is read from shared memory. A new implementation making use of disabled Global Interpreter Lock (GIL) further reduces runtime. The runtime generally scales linearly with the amount of cells simulated. Scaling with longer simulated time is less than linear, as cells in compartment $C_{2}$ are removed from the simulation after storing their location. The process of data preparation on the other hand also depends on specified parameters; here the resolution of pipe volumes that vessels are separated into, voxel sizes and desired distance between reference nodes inside of vessel meshes play a crucial role. In this step the flow fields are also calculated using a CUDA kernel. Once the data preparation is finished, it is stored and can be loaded for consecutive simulations. Our simulation of 100,000 cells over a duration of 3600 seconds took about 45 minutes on an AMD EPYC 8324P 32-Core Processor (64 threads) using 32 spawned processes. The data preparation beforehand took approximately 1.5 hours.
@@
The here presented simulation framework, CeFloPS, has been demonstrated to be capable of simulating physiologically meaningful cell pathways within the human body, which can be utilized for the development of image reconstruction algorithms as well as specific cell tracking strategies. CeFloPS allows to simulate not only the flow of cells through the vessel branches but also the migration of cells from the vasculature into tissues either as reversible or non-reversible trapping steps. By that, different physiologically relevant scenarios can be simulated for various cell types providing viable data for mimicking the dynamic distribution of cells within the human body. Beyond pure cell-based simulations, even tracer studies could be simulated by CeFloPS, as shown here by the FDG example. Following the CeFloPS simulations, a label can be assigned to the cells, which was chosen for PET studies as a radio-label ($\beta^{+}$ emitter) but could be of any kind allowing also simulations for other imaging modalities, such as SPECT, OI, MRI, or ultrasound imaging.
The cell flow simulation presented is demonstrated and partially validated using FDG as a model for ”cell” motion. As FDG is a well known tracer for PET which provides reference values from numerous publications, organ-specific compartment model parameters and resulting tracer concentrations could be utilized for our study. Such information is currently hardly available from cell tracking studies making a translation towards specific cell types (e.g. T-cells, immune cells, etc.) difficult to develop and validate. As specific data is available for specific cell flow behavior in humans, dedicated CeFloPS simulations can be developed. Comparing the MIP of the cell distribution in the first 100 seconds to similar MIPs of dynamic total-body FDG PET scans like [42] and [43], a similar flow behavior can be observed.
@@
The radioactivity of 1 kBq attributed to each FDG “cell” in the example is quite high and not realistic in the context of cell tracking experiments. In any case, this served only as a proof-of-principle value allowing a relatively noise-free dynamic image reconstruction of the simulated PET data for comparison with ground-truth cell positions. For realistic cell tracking simulation, radioactivity values in the range of a few Bq would be more appropriate. However, the demonstration of cell trajectory based PET reconstructions for low-count data, as shown before by our group (see [22]), was not part of this work and will be subject of subsequent studies.
@@
It is evident from the cell numbers (Fig. 9) that certain organs are preferentially traversed by cells over others. This is due to the fact that the probability of which blood vessel branch is chosen at bifurcations is dependent upon the proportion of blood flow, as illustrated by (5). Organs that are less likely to be traversed by cells will therefore have a lower cell count. This is due to the rudimentary modeling of the blood vessels in the kidneys by XCAT, exhibiting a lack of small diameter vessels (see Fig. 1c). Given that these represent a significant resistance, as indicated by (2), the blood flow to the kidneys is currently overestimated. Consequently, the likelihood of cell paths traversing the kidneys is greater than that of other organs. The spatial resolution of the XCAT phantom thus represents a clear limitation of CeFloPS. In order to improve towards a more realistic behavior, two potential approaches could be followed. Either, missing smaller blood vessels could be added to the XCAT phantom which would lead to a higher resistance in the vasculature and thus to a more realistic blood flow. A potential approach is illustrated in [44], which outlines techniques for extending the pulmonary vasculature in the lungs. Or, pre-defined resistances could be attributed to dedicated organs which would also lead to a more realistic flow simulation in CeFloPS. In principle, a different anthropomorphic phantom with higher level of details could also be implemented in the CeFloPS framework which could circumvent the flow imbalances.
@@
Blood flow is currently implemented as a continuous laminar flow, completely ignoring the pulsating characteristics of the blood circulation. The assumption of laminar flow in the blood vessels is a highly simplified model that allows for rapid determination of cell positions within the vessels without great computational effort and is supported by the Reynolds numbers in the blood vessels, which were obtained for all vessels. Reynolds numbers are all well below 2000, which is a known threshold after which laminar flow slowly transitions into turbulent flow (Reynolds number > 4000) [45]. For example, the Reynolds numbers of the aorta range from 1300 to 1900. In addition, no pulsative blood flow is allowed in this state of the simulation which further supports our assumption. A more comprehensive depiction of blood flow would encompass pulsed flow, along with additional parameters such as stress in arterial walls or hemorheology and would require solving the Navier-Stokes equation, as outlined for example in [46] and [47]. Other approaches involve the utilization of finite element or Lattice Boltzmann methods to model blood flow within the vessels, the latter having previously been employed for the arterial vessels of the XCAT phantom, as demonstrated in [15]. With regard to the intended field of application of CeFloPS, namely the generation of anatomically and physiologically meaningful cell pathways for PET simulations, the computational effort of these methods exceeds the purpose of our study. Nevertheless, these methods could be of interest when focusing on pulsed flow and local effects due to the pulse wave behavior. Also, turbulent flow and a variable blood viscosity which is known to play an important role in the area of the capillary bed, are completely ignored. Moreover, cells in our study are considered as small particles which freely flow through the blood system and freely diffuse through the tissue compartments. Cell specific behaviors, such as rolling or adhesion at the endothelium or immune related activation cannot explicitly be modeled within the CeFloPS simulation. However, a similar behavior of cells could potentially be modeled by setting (dynamic) $k$-Parameters (Tab. LABEL:tab:k_values) for compartmental exchanges.

### v3

@@ 2.4 PET simulation
Raw PET data are generated using the Geant4 application for tomographic emission GATE (version 9.0) [41, 19]. The cell path outputs from the CeFloPS simulations have the format of GATE placement files and thus are used as generic moving sources in the simulations. The sources, i.e. the cells, are simulated as spheres whose diameters and activities can be specified by the user. For the example shown in this work, a diameter of 10 µm and an activity of 1 kBq per cell were simulated for 100 cells. The activity per cell and number of cells do not reflect a realistic scenario but were chosen solely as a proof-of-principle. The PET scanner used in the simulation is based on the Siemens Biograph Vision Quadra (Siemens Healthineers, Erlangen, Germany), a large-body PET system using approximately 250000 single lutetium oxyorthosilicate (LSO) detector elements covering an axial field of view (FOV) of 106 cm [42]. The background radiation in the scanner generated by the presence of 176Lu in the LSO crystals can be simulated if desired but was neglected in the current simulation. The attenuation map for gamma energies of 511 keV generated from the XCAT phantom was used as an attenuation phantom. The attenuation phantom was positioned in the PET scanner so that the upper body, including the head, was within the FOV of the scanner. Since the simulated cell paths are based on the blood vessel and organ structure of the XCAT phantom, their positions correspond exactly to the anatomical structures in the attenuation phantom, making the simulation also realistic in view of the proportion of attenuated and scattered events. Although the lower body extends beyond the scanner’s FOV, it is still simulated to account for cells located there. The output of the GATE simulations is provided by ROOT files, which can be converted into a format required for the PEPT or PET algorithm applied by the user, e.g. list mode format or sinograms. In this work, no dedicated cell tracking algorithm was applied. Instead, the ROOT output was converted into list mode format and was reconstructed using a standard expectation-maximization algorithm to demonstrate that the simulations yield realistic data.
@@
Starting from the aorta, with a mean blood flow of 100100~ml/s, the blood distributes into all vessels and tissues depending on the resistances of the vessel system and the blood pressure differences (as listed exemplarily in Table 2). It is evident that the lungs receive the entire cardiac output through the right ventricular contraction, which accounts for 100 % of the total blood flow. Notably, the kidneys, gastrointestinal tract, and liver show strong alignment between the simulated flow rate and physiological flow distributions from literature [43], highlighting CeFLoPS’s ability to capture predominant systemic blood flow patterns. However, considerable discrepancies are observed for the brain, spleen, pancreas, and thyroid. This behavior can be attributed to the insufficient definition of the vascular structure of these organs or a high vessel resistance due to a large number of small vessels in the case of the brain.
Table 2: Blood flow distribution simulated with CeFloPS
@@
(% of total blood flowa)
Reference blood flow rate
in humansb
(% of total blood flow)
@@
Table 3 presents the mean blood flow rate and Reynolds number for the blood vessels in CeFloPS. Vessels are grouped into three size-based categories separately for arteries and veins. As expected from (1) to (5), large arteries such as the aorta exhibit the highest flow rates and Reynolds numbers, whereas small arteries and veins show markedly lower values, consistent with their higher resistance.
Table 3: Average vessel radius ($\overline{r}$), blood flow rate ($\overline{Q}$), and Reynolds number ($\overline{\mathrm{Re}}$) for different vessel types in CeFloPS, differentiated between arteries and veins and further divided into large ($r\geq 6\,\mathrm{mm}$), medium ($3\,\mathrm{mm}\leq r<6\,\mathrm{mm}$), and small ($r<3\,\mathrm{mm}$) groups
Vessel type $\overline{r}$ [$m$] $\overline{Q}$ [ml/s] $\overline{\mathrm{Re}}$
large arteries (e.g. aorta) 8.49 70.69 1568
medium arteries (e.g. arteriae femorales) 3.97 16.03 685
small arteries (e.g. arteriae renales) 0.81 0.39 58
large veins (e.g. venae cavae) 6.51 35.12 1041
medium veins (e.g. venae hepaticae) 3.92 7.24 357
small veins (e.g. sinus sigmoidei) 0.96 0.59 82
Fig. 6 depicts a maximum intensity projection (MIP) of the distribution of 100,000 injected cells over the first 100 seconds following injection into the left arm vein. As would be anticipated, the route from the injection site to the heart, the heart itself, the pulmonary arteries and veins, the aorta and arteries close to the aorta are traversed with high frequency. In regions such as the legs and arms, it is evident that certain arteries and veins have yet to be reached by a significant number of cells. Additionally, there are observable transitions of cells into organs and backflow into veins, particularly evident in the kidneys, liver, and vena cava.
Figure 6: Maximum intensity projection (1.5 mm pixel size) of the vessels and tissues traversed by 100,000 cells injected into the right arm vein during the first 100 seconds of a simulation. The gray log scale encodes the number of cells per mm2.
The biased 3D random walk that cells perform when they migrate from the blood into a tissue is illustrated in Fig. 7a using the example of the lung for a duration of 2000 s. The simulation outcome is visualized through MIPs in three orthogonal planes to clearly illustrate spatial traversal patterns. The distribution of cells entering the random walk through the arterial sources is weighted by simulated arterial blood flow fractions of the vessels in the lung. It prominently demonstrates regions of enhanced cellular traversal intensity, indicative of high-flow vascular pathways, and shows that the biased random walk ensures that the entire organ tissue is potentially accessible by incoming cells. Furthermore, the temporal dynamics of cell migration are elucidated in Fig. 7b, where the histogram of random walk durations reveals a rapid traversal of the lung vasculature for the majority of cells. The mean duration is 3.77 s, consistent with the theoretical specification outlined in [32].
@@
Fig. 8 illustrates the locations of 100 simulated cells within a node system at six time points, corresponding to the output of the flow model, compared to MIPs reconstructed from PET data using an expectation-maximization based image reconstruction without attenuation correction. The reconstructed positions and cell counts align well with ground-truth positions, despite minor reconstruction artifacts causing varying intensities in the MIPs due to missing attenuation and scatter correction. Initially, cells (red spheres) are injected into the left arm vein (median cubital) over ten seconds. After six seconds, the first cells migrate up the vein while new cells are still injected. By 18 to 30 seconds, cells have reached the left subclavian vein and heart and start to move into the lungs. At 60 seconds, the majority of cells are located in vasculature and organs of the thorax, with visible spread towards the lungs and kidneys. A slowly traveling cell remaining near the injection site can be observed. Over longer timescales (300–600 s), the distribution extends across the torso, with cells accumulating in the organs and transitioning to compartments $C_{1}$ or $C_{2}$, changing color from red to green. In the MIP, this leads to spots of higher intensity, most notably visible in the kidneys, thyroids, and the brain.
@@
Figure 8: Output of the flow model (left) showing simulated ground-truth cell positions for 100 cells at six different time points, including both moving cells (red) in the node system as well as stationary cells that have entered $C_{1}$ or $C_{2}$ (green). For illustration, PET maximum intensity projections (right) using a standard EM based image reconstruction show the cell distribution (black) superimposed on the organ and blood vessel structure. The red dotted line marks the FOV of the simulated PET scanner.
@@
Figs. 9 and 10 illustrate the temporal distribution of 100,000 cells in arterial blood, lung, kidneys, liver, spleen, and pancreas as absolute cell numbers (sum of cells in $C_{1}$ and $C_{2}$) and as cell concentrations per milliliter tissue, respectively. Initially, the cell count in the arterial blood rapidly drops and a pronounced peak of cell presence is observed in the lung, liver, and kidneys, indicative of early vascular dissemination and transitions from $C_{A}$ to $C_{1}$. Subsequently, cell counts decline within these tissues, aligning with physiological expectations of cells transitioning back from $C_{1}$ to $C_{A}$ and leaving the tissue into the venous blood. Notably, over extended simulation time, the cell count in the kidneys rises again and the kidneys and the pancreas exhibit a progressive increase in cell accumulation, indicating trapping of cells in $C_{2}$. Meanwhile, the spleen maintains a relatively steady cell population, reflecting a possible equilibrium between cell influx and clearance. The concentration plots (Fig. 10) underscore these observations by correcting for organ volume, thus offering additional insight into organ-specific cell affinity beyond raw cell counts. In terms of raw cell count, spleen and pancreas reach significantly lower numbers than the other tissues, which is not apparent in the concentration plot. Notably, the kidneys exhibit considerably higher values than other tissues with regard to both the absolute number of cells and the cell concentration.
@@
Fig. 11 illustrates the temporal evolution of the 100,000 simulated cells across the virtual, blood, and tissue compartments depicted in Fig. 4 and described by (12). After the 30 s injection period, all cells have transitioned from $C_{\mathrm{inj}}$ to $C_{A}$. This is followed by a swift redistribution into tissues or removal from blood and transition into $C_{\mathrm{rem}}$. After a brief transient phase, a stable equilibrium emerges, characterized predominantly with over 80% of cells in the removed state. The dynamics clearly show that only a small fraction of injected cells persist within blood or the defined tissue compartments at longer time points. The majority of cells are located in tissues without specified dynamics, like skin, muscle or adipose.
@@
Figure 11: Temporal distribution of 100,000 simulated cells across compartments shown in Fig. 4: injected, arterial blood, tissue, and removed.
@@
Fig. 12 compares the simulated cell concentrations in the tissues to the distributions one would expect from (11). To calculate the theoretically expected distributions for each tissue, the number of cells perfusing from the arterial blood into tissue are mapped. The arterial input to each tissue is determined by the unique and locally dependent distribution of blood flow and the subsequent delay and dispersion effects. Together with the transition rates (Table 1) of each tissue, the arterial input allows derivation of the expected cell distributions from the differential equations. The reference volume to which the cells of the arterial input function are normalized in order to obtain a concentration is treated as a tissue-specific variable. This variable is dependent on the volume of the small arteries, arterioles and capillaries leading to and perfusing the tissue. This factor was determined by fitting the two distributions and multiplying the result with the arterial input function. This yields the concentration of cells in the arterial blood and the expected theoretical curve. As shown in fig. 12 the simulated distributions describe the theoretically expected ones well, with small over- or undershoots of the initial peaks. In the case of the spleen and pancreas, the theoretical expected cell concentration exhibits fluctuations, due to the small number of cells in these tissues.
@@
Figure 12: Simulated concentration distributions profiles of cells in tissues and tissue specific arterial blood input. Solid colored lines represent the simulated cell concentrations per milliliter in the respective tissues, whereas the red lines show the cell concentration in the arterial blood perfusing each tissue. The plot in the upper left displays the total arterial blood concentration in the whole body. The dashed black lines depict the theoretically expected concentration distributions derived from (11), based on the tissue specific arterial input functions and transition rates (Table 1).
@@
The simulation framework presented here, CeFloPS, has been demonstrated to be capable of simulating physiologically meaningful cell pathways within the human body, which can be utilized for the development of image reconstruction algorithms as well as specific cell tracking strategies. CeFloPS allows to simulate not only the flow of cells through the vessel branches but also the migration of cells from the vasculature into tissues either as reversible or non-reversible trapping steps. In this way, different physiologically relevant scenarios can be simulated for various cell types providing viable data for mimicking the dynamic distribution of cells within the human body. Beyond pure cell-based simulations, even tracer studies could be simulated by CeFloPS, as shown here by the FDG example. Following the CeFloPS simulations, a label can be assigned to the cells, which was chosen for PET studies as a radio-label ($\beta^{+}$ emitter) but could be of any kind allowing also simulations for other imaging modalities, such as SPECT, OI, MRI, or ultrasound imaging.
The cell flow simulation presented is demonstrated and partially validated using FDG as a model for ”cell” motion. As FDG is a well known tracer for PET which provides reference values from numerous publications, organ-specific compartment model parameters and resulting tracer concentrations could be utilized for our study. Such information is currently hardly available from cell tracking studies making translation towards specific cell types (e.g. T-cells, immune cells, etc.) difficult to develop and validate. As specific data is available for specific cell flow behavior in humans, dedicated CeFloPS simulations can be developed. Comparing the MIP of the cell distribution in the first 100 seconds (Fig. 6) and the simulated position of 100 cells over 600 seconds (Fig. 8) to similar MIPs of dynamic total-body FDG PET scans like [44] and [45], a similar flow behavior can be observed. Both the cells in the simulations and the FDG in the PET scans take approximately 20 to 30 seconds to traverse the distance from the injection site (arm or leg, respectively) to the heart and lungs. The PET scans show FDG accumulation in the kidneys after 35 to 40 seconds, which subsequently maintain a high signal over the entire scan time. In the simulations, the cell accumulation in the kidneys is observable after approximately 60 seconds and persists until the maximum simulation time of 10 minutes. Large blood vessels can still be distinctly identified in the scans after 10 minutes due to their signal intensity. This is also consistent with the simulations, where after 10 minutes a significant proportion of the cells are located in the vascular system, as can be seen from the color coding in Fig. 8.
@@
The radioactivity of 1 kBq attributed to each FDG “cell” in the example is quite high and not realistic in the context of cell tracking experiments. In any case, this served only as a proof-of-principle value allowing a relatively noise-free dynamic image reconstruction of the simulated PET data for comparison with ground-truth cell positions. For realistic cell tracking simulation, radioactivity values in the range of a few Bq would be more appropriate. However, the demonstration of cell trajectory based PET reconstructions for low-count data, as shown before by our group (see [23]), was not part of this work and will be the subject of subsequent studies.
@@
It is evident from the cell numbers (Fig. 9) that certain organs are preferentially traversed by cells over others. This is due to the fact that the probability of which blood vessel branch is chosen at bifurcations is dependent upon the proportion of blood flow, as illustrated by (6). Organs that are less likely to be traversed by cells will therefore have a lower cell count. This is due to the rudimentary modeling of the blood vessels in the kidneys in XCAT, exhibiting a lack of small diameter vessels (see Fig. 1c). Given that these represent a significant resistance, as indicated by (2), the blood flow to the kidneys is currently overestimated. Consequently, the probability of cell paths traversing the kidneys is greater than that of other organs. The spatial resolution of the XCAT phantom thus represents a clear limitation of CeFloPS. In order to improve towards a more realistic behavior, two potential approaches could be followed. Either, missing smaller blood vessels could be added to the XCAT phantom which would lead to a higher resistance in the vasculature and thus to a more realistic blood flow. A potential approach is illustrated in [46], which outlines techniques for extending the pulmonary vasculature in the lungs. Or, pre-defined resistances could be attributed to dedicated organs which would also lead to a more realistic flow simulation in CeFloPS. In principle, a different anthropomorphic phantom with higher level of details could also be implemented in the CeFloPS framework which could circumvent the flow imbalances.
@@
Blood flow is currently implemented as a continuous laminar flow, completely ignoring the pulsating characteristics of the blood circulation. The assumption of laminar flow in the blood vessels is a highly simplified model that allows for rapid determination of cell positions within the vessels without great computational effort and is supported by the Reynolds numbers in the blood vessels, which were obtained for all vessels. All Reynolds numbers are well below the threshold value of 2000 at which transition to turbulent flow would occur [30]. For example, the Reynolds numbers of the aorta range from 1300 to 1900. In addition, no pulsatile blood flow is allowed in this state of the simulation which further supports our assumption. A more comprehensive depiction of blood flow would encompass pulsed flow, along with additional parameters such as stress in arterial walls or hemorheology and would require solving the Navier-Stokes equation, as outlined for example in [47] and [48]. Other approaches involve the utilization of finite element or Lattice Boltzmann methods to model blood flow within the vessels, the latter having previously been employed for the arterial vessels of the XCAT phantom, as demonstrated in [16]. With regard to the intended field of application of CeFloPS, namely the generation of anatomically and physiologically meaningful cell pathways for PET simulations, the computational effort of these methods exceeds the scope of our study. Nevertheless, these methods could be of interest when focusing on pulsed flow and local effects due to the pulse wave behavior. Also, turbulent flow and a variable blood viscosity, which is known to play an important role in the area of the capillary bed, are completely ignored. Moreover, cells in our study are considered as small particles, which freely flow through the blood system and freely diffuse through the tissue compartments. Cell-specific behaviors, such as rolling or adhesion at the endothelium or immune related activation cannot explicitly be modeled within the CeFloPS simulation. However, a similar behavior could potentially be modeled by setting (dynamic) $k$-parameters (Tab. 1) for compartmental exchanges.